A cargo plane flew from the US across the Atlantic at two hundred forty three mph and flew back to the US at two hundred sixty three mph. Given that the first trip took two hours long was the flight to NYC.
Answer by Boreal(15235)

speed*time=distance
243*t=243t
back
263(t-2)=263t-526
the distances are equal
243t=263t-526
-20t=-526
t=26.3 hours over
t-2=24.3 hours back ANSWER
Multiply that by the speed and the distance is 6390.9 miles.